# A Liquid-Core Fiber Platform for Classical and Entangled Two-Photon Absorption Measurements

### v1

@@ 9 Calculating a C2PA Cross Section
The uncertainty on derived C2PA cross sections was determined by propagating the errors of the various parameters that go into Eq. (9). We multiply this value by a coverage factor ($k=2$) to reach $\approx$ 95% confidence that the true value lies within the bounds set by the error bars. The uncertainty from each experiment is 34%.
@@
In Figure 8, we plot concentration normalized C2PEF ($\frac{F_{C}}{c}$) at $100$ nW from measurements (fit) and calculations (calc) for the experimental conditions of experiments 1 (blue), 2 (orange) and 3 (green). The calculations use Eq. (18) with the values of parameters input from Tables 2 and 3. The uncertainty on these values is propagated, and multiplied by a coverage factor ($k=2$) to give a total uncertainty of 38% on the calculation (shown by shaded regions). After each experiment, small improvements to the setup were made that resulted in a shift of the curve upwards on the plot as experiments progressed from 1 to 3. The trend of each curve shows the same decrease of $\frac{F_{C}}{c}$ as a function of $c$. This general decrease is due to fluorescence reabsorption by the abundant molecules available at high concentrations. The data points from measurements are extracted using quadratic fits (slope fixed to 2.00) to each experiments’ data of fluorescence as a function of power. The fit is extrapolated to $100$ nW and that value is divided by the measured concentration of the experiment’s sample. These data points are within the expected range based on the calculations and their respective uncertainties.
10 Calculating an E2PA Cross Section Upper Bound
In this section, in a similar manner to Section 9, we discuss the assumptions and equations used to calculate an E2PEF signal. We arrive at an equation to derive an E2PA cross section upper bound.
@@ 10 Calculating an E2PA Cross Section Upper Bound
where $N_{E}(z)$ (excitations cm-1 pulse-1) is the number of excitations per infinitesimal length of fiber $dz$ (cm) per pulse and the other parameters are as mentioned in Section 9. The parameter $N_{E}(z)$ is defined as
$\begin{split}N_{E}(z)&=\frac{1}{2}\sigma_{E}(z)n\mathcal{T}\eta_{T}(\lambda_{e},z)\frac{Q(z)}{g}\\ &=\frac{1}{2}\sigma_{E}(z)n\mathcal{T}\eta_{C}\eta_{A}(\lambda_{e},z)\eta_{S}(\lambda_{e},z)\frac{Q(z)}{g},\end{split}$ (21)
where $\sigma_{E}(z)$ (cm2 fluorophore-1) is the E2PA cross section, which has $z$ dependence due to group delay dispersion broadening the entanglement time, $\mathcal{T}$ is the free-space transmission efficiency of the SPDC from the center of the crystal to the fiber, and $\eta_{T}$ and $\eta_{C}$ are the fiber transmission and coupling efficiencies (Section 6). The other parameters are as mentioned in Section 9.
The parameters $\mathcal{T}$ and $\eta_{T}$ are required for E2PA calculations but not for C2PA calculations because E2PA requires that photon pairs stay intact whereas C2PA only requires the presence of single photons. This can be understood by writing the SPDC pair photon rate $Q_{\mathrm{pair}}(z)$ (photon pair s-1) at the fiber as
$\begin{split}Q_{\mathrm{pair}}(z)&=\frac{1}{2}\mathcal{T}^{2}\eta_{T}^{2}(\lambda_{e},z)Q_{\mathrm{xtal}}\\ &=\frac{1}{2}\mathcal{T}\eta_{T}(\lambda_{e},z)Q(z),\end{split}$ (22)
where $Q_{\mathrm{xtal}}$ (photon s-1) is the (single) photon rate at the crystal and $Q(z)$ is the (single) photon rate at the fiber. More details on this can be found in Ref. 16.
Here we will use the probabilistic model (see for example Ref. 16) to describe $\sigma_{E}(z)$ (cm2 molecule-1) as
@@
We simulate the entanglement time using a discrete Fourier transform of the measured SPDC spectrum shown in Ref. 16. To do this, we estimate the joint spectral amplitude as
@@
where $F(\omega_{\mathrm{S}},\omega_{\mathrm{I}})$ is the measured joint spectral intensity, and $\omega_{\mathrm{S}}$, $\omega_{\mathrm{I}}$ and $\omega_{\mathrm{P}}$ are the signal, idler and pump frequencies and the parameters $D_{0}$ and $\beta$ are described in Section 9. Discrete Fourier transforms are performed at discrete steps along the fiber of 1 cm and the entanglement time is calculated as the FWHM of projection of the joint temporal intensity along the ($t_{\mathrm{S}}-t_{\mathrm{I}}$) axis. Here $t_{\mathrm{S}}$ and $t_{\mathrm{I}}$ are the time of arrival of signal and idler respectively. The FWHM is calculated from standard deviation due to the spectrum’s non-Gaussian nature. We fit the $z$-dependence of the resulting $T_{e}$ values using
@@
In order to estimate a single value for the E2PA cross section, we will use
@@
$N_{E}(z)=\frac{1}{2}\sigma_{E}(0)n\frac{T_{e}(0)}{T_{e}(z)}\mathcal{T}\eta_{T}(\lambda_{e},z)\frac{Q(z)}{g}.$ (27)
@@
$F_{E}=\frac{1}{2}\sigma_{E}(0)n\mathcal{T}\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}\eta_{T}(\lambda_{e},z)Q(z)$
@@
$\sigma_{E}(0)=2F_{E}\times\left(n\mathcal{T}\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}\eta_{T}(\lambda_{e},z)Q(z)\right.$
@@
Then for a null measurement, we replace $F_{E}$ by the measurable fluorescence lower bound $F^{\mathrm{LB}}$ (cnt s-1) and $\sigma_{E}(0)$ becomes $\sigma_{E}^{UB}$,
$\sigma_{E}^{UB}=2F^{\mathrm{LB}}\times\left(n\mathcal{T}\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}\eta_{T}(\lambda_{e},z)Q(z)\right.$
@@
All the parameters in Eq. (10) are known through experiments or simulations. The parameter $F^{\mathrm{LB}}$ is estimated to be $1.0$ cnt s-1 based on the signal levels that could be distinguished in these measurements, $\mathcal{T}$ is measured, $T_{e,0}$ and $S_{0}$ are estimated using the discrete Fourier transform of the measured joint spectrum, $D_{0}$ is estimated, $Q(z)$ is measured, and all other parameters are found as stated in Section 9. The values of these parameters are listed in Tables 2 and 3.
The uncertainty on the derived E2PA cross section upper bound was determined by propagating the errors of the various parameters that go into Eq. (10) and multiplying this value by a coverage factor ($k=2$), which gives 40%.

### v2

@@ Conclusions
We can solve for $\phi_{0}(z)$ by inputting Eq. (Calculating a C2PA Cross-Section) into Eq. (17) and integrating over $x$, $y$ and $t$,
Now we can rewrite Eq. (11) using these equations as
[…]
where $\lambda_{e}=810$ nm.
Then we can solve for the C2PA cross-section,
All the parameters in Eq. (Calculating a C2PA Cross-Section) are known through experiments or simulations. The parameters $d_{0}$ and $\beta$ are estimated using COMSOL Multiphysics Simulation software, $F_{C}/W_{0}^{2}$ (cnt s-1 $\mu$W-2) is the fit to our experimental C2PEF data (and $W_{0}$ is estimated using measured $W(l)$), $n$ is measured using a spectrophotometer, $a_{\mathrm{sol}}(\lambda)$ is from literature40, $\mu(\lambda)$ is determined from scattering measurements, $\tau_{0}$ is measured using a SwampOptics Grenouille 8-50-USB, $D_{0}$ is measured, $k(\lambda)$ is calculated, $\Phi(\lambda)$ is known from published AF455 measurements47, 48 and $\gamma(z,\lambda)$ is calculated based on optics’ specifications, our scattering measurements and the published extinction coefficient and spectra of AF45548. Table 2 and Table 3 summarize parameter values used in this calculation and that described in the following section.
The uncertainty on derived C2PA cross-sections was determined by propagating the errors of the various parameters that go into Eq. (Calculating a C2PA Cross-Section). We multiply this value by a coverage factor ($k=2$) to reach $\approx$ 95% confidence that the true value lies within the bounds set by the error bars. The uncertainty from each experiment is 34%.
@@
In Figure 8, we plot concentration normalized C2PEF ($\frac{F_{C}}{c}$) at $100$ nW from measurements (fit) and calculations (calc) for the experimental conditions of experiments 1 (blue), 2 (orange) and 3 (green). The calculations use Eq. (20) with the values of parameters input from Tables 2 and 3. The uncertainty on these values is propagated, and multiplied by a coverage factor ($k=2$) to give a total uncertainty of 38% on the calculation (shown by shaded regions). After each experiment, small improvements to the setup were made that resulted in a shift of the curve upwards on the plot as experiments progressed from 1 to 3. The trend of each curve shows the same decrease of $\frac{F_{C}}{c}$ as a function of $c$. This general decrease is due to fluorescence reabsorption by the abundant molecules available at high concentrations. The data points from measurements are extracted using quadratic fits (slope fixed to 2.00) to each experiments’ dataset of fluorescence as a function of power. The fit is extrapolated to $100$ nW and that value is divided by the measured concentration of the experiment’s sample. These data points are within the expected range based on the calculations and their respective uncertainties.
Calculating an E2PA Cross-Section Upper Bound
In this section, in a similar manner to the previous section, we discuss the assumptions and equations used to calculate an E2PEF signal. We arrive at an equation to derive an E2PA cross-section upper bound.
@@
where $N_{E}(z)$ (excitations cm-1 pulse-1) is the number of excitations per infinitesimal length of fiber $dz$ (cm) per pulse and the other parameters are as mentioned in the previous section. The parameter $N_{E}(z)$ is defined as
$N_{E}(z)=\sigma_{E}(z)n\frac{Q_{\mathrm{pairs}}(z)}{g}$ (23)
where $\sigma_{E}(z)$ (cm2 fluorophore-1) is the E2PA cross-section, which has $z$ dependence due to group delay dispersion broadening the entanglement time and $Q_{\mathrm{pairs}}(z)$ (photon pairs s-1) is the SPDC photon pair rate in fiber. The other parameters are as mentioned in the previous section.
In contrast to C2PA, E2PA requires the presence of spatially and temporally correlated photon pairs. Furthermore the excitation rate scales linearly, as opposed to quadratically, with photon rate in the low-gain regime. In this section, we use the scaling in the low-gain regime to make the most conservative estimate of the upper bound of the E2PA cross-section. In order to write an equation for $Q_{\mathrm{pairs}}(z)$, we will first write the SPDC single photon rate at the fiber, $Q(z)$ (photons s-1),
$Q(z)$ $=$ $\eta_{F}\eta_{C}\eta_{A}(\lambda_{e},z)\eta_{S}(\lambda_{e},z)\frac{Q^{\mathrm{mm}}_{\mathrm{xtal}}}{M}$
$=$ $\eta_{F}\eta_{C}\eta_{A}(\lambda_{e},z)\eta_{S}(\lambda_{e},z)Q_{\mathrm{xtal}},$
where $Q^{\mathrm{mm}}_{\mathrm{xtal}}$ (photons s-1) is the total (all modes) single photon rate at the crystal, $M$ is the number of spatial modes of the generated SPDC, $Q_{\mathrm{xtal}}$ (photons s-1) is the single photon rate at the crystal for the spatial mode coupled in fiber, and the other parameters are as mentioned in previous sections.
We can then write the SPDC photon pair rate at the fiber in terms of the single photon rate,
$Q_{\mathrm{pairs}}(z)$ $=$ $\eta_{K}\eta_{A}(\lambda_{e},z)\eta_{S}(\lambda_{e},z)\frac{Q(z)}{2}$
$=$ $\eta_{K}\eta_{F}\eta_{C}\eta_{A}^{2}(\lambda_{e},z)\eta_{S}^{2}(\lambda_{e},z)\frac{Q_{\mathrm{xtal}}(z)}{2}$
$=$ $\eta^{{}^{\prime}}_{K}\eta_{F}^{2}\eta_{C}^{2}\eta_{A}^{2}(\lambda_{e},z)\eta_{S}^{2}(\lambda_{e},z)\frac{Q_{\mathrm{xtal}}(z)}{2},$
where $\eta_{K}^{{}^{\prime}}$ is the effective Klyshko efficiency and $\eta_{K}$ is the Klyshko efficiency as defined in Eq. (Experimental Characterization). The third line of this equation illuminates a signature of E2PA—while E2PA scales linearly with photon rate, it scales quadratically with single photon loss. More details on this signature can be found in Ref. 16.
Here we will use the probabilistic model (see for example Ref. 16) to describe $\sigma_{E}(z)$ as
@@
We simulate the entanglement time using a discrete Fourier transform of the measured SPDC joint spectrum shown in Ref. 16. To do this, we estimate the joint spectral amplitude as
@@
where $F(\omega_{\mathrm{S}},\omega_{\mathrm{I}})$ is the measured joint spectral intensity, and $\omega_{\mathrm{S}}$, $\omega_{\mathrm{I}}$ and $\omega_{\mathrm{P}}$ are the signal, idler and pump frequencies and the parameters $D_{0}$ and $\beta$ are described in the previous section. Discrete Fourier transforms are performed at discrete steps along the fiber of 1 cm and the entanglement time is calculated as the FWHM of the projection of the joint temporal intensity along the ($t_{\mathrm{S}}-t_{\mathrm{I}}$) axis. Here $t_{\mathrm{S}}$ and $t_{\mathrm{I}}$ are the time of arrival of signal and idler respectively. The FWHM is calculated from standard deviation due to the spectrum’s non-Gaussian nature. We fit the $z$-dependence of the resulting $T_{e}$ values using
@@
In order to estimate a single value for the E2PA cross-section, we will use
@@
$N_{E}(z)=\sigma_{E}(0)n\frac{T_{e}(0)}{T_{e}(z)}\frac{Q_{\mathrm{pairs}}(z)}{g}.$ (30)
@@
$F_{E}=\sigma_{E}(0)n\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}Q_{\mathrm{pairs}}(z)$
@@
$\sigma_{E}(0)=F_{E}\times\left(n\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}Q_{\mathrm{pairs}}(z)\right.$
@@
Then for a null measurement, we replace $F_{E}$ by the measurable fluorescence lower bound $F^{\mathrm{LB}}$ (cnt s-1) and $\sigma_{E}(0)$ becomes $\sigma_{E}^{\mathrm{UB}}$,
$\sigma_{E}^{\mathrm{UB}}=F^{\mathrm{LB}}\times\left(n\int_{0}^{l}\frac{T_{e}(0)}{T_{e}(z)}Q_{\mathrm{pairs}}(z)\right.$
@@
All the parameters in Eq. (Calculating an E2PA Cross-Section Upper Bound) are known through experiments or simulations. The parameter $F^{\mathrm{LB}}$ is estimated to be $1.0$ cnt s-1 based on the signal levels that could be distinguished in these measurements, $T_{e,0}$ and $S_{0}$ are estimated using the discrete Fourier transform of the measured joint spectrum, $D_{0}$ is estimated, $Q_{\mathrm{pairs}}(z)$ is calculated using the measured value of $Q_{\mathrm{out}}$ and measured and estimated values of efficiencies, and all other parameters are found as stated in the previous section. The values of these parameters are listed in Tables 2 and 3.
The uncertainty on the derived E2PA cross-section upper bound was determined by propagating the errors of the various parameters that go into Eq. (Calculating an E2PA Cross-Section Upper Bound) and multiplying this value by a coverage factor ($k=2$), which gives 40%.